# Efficiently learning depth-3 circuits via quantum agnostic boosting

### v1

@@ Abstract
    $s$-term DNF formulas in near-polynomial time $\textsf{poly}(n,(s/\varepsilon)^{\log\log s/\varepsilon})$.
@@
Using quantum agnostic boosting, we obtain the first “near” polynomial time $n^{O(\log\log n)}$ algorithm for learning $\textsf{poly}(n)$-sized depth-$3$ circuits (consisting of AND, OR, NOT gates) in the uniform quantum PAC model using quantum examples. Classically, the analogue of efficient learning depth-$3$ circuits (and even depth-$2$ circuits) in the uniform PAC model has been a longstanding open question in computational learning theory. Our work nearly settles this question, when the learner is given quantum examples.
@@ 1.1 Main results
Our main result is to answer both of the questions highlighted above in the affirmative. For both the tasks of agnostically learning $\operatorname{poly}(n)$-sized DNFs and PAC learning $\operatorname{poly}(n)$-sized depth-$3$ circuits, we give $n^{O(\log\log n)}$ time quantum algorithms. Below, we will discuss these results, starting with our main technical contribution, quantum agnostic boosting.
@@
Size-$t$ decision trees and $k$-juntas are learnable in time $\operatorname{poly}(n,t,1/\varepsilon)$ and $\operatorname{poly}(n,2^{k},1/\varepsilon)$ respectively. Similarly, $s$-term DNF formulas are learnable in time $\operatorname{poly}(n,(s/\varepsilon)^{\log\log s/\varepsilon})$.
We formally state this theorem for each class in Section 4. In order to compare with classical results and for ease of exposition, let us consider the parameters above to take values of $2^{k},t,s,1/\varepsilon=\operatorname{poly}(n)$ (which in general are considered the most interesting settings of these parameters for learning). Classically, the SOTA algorithm for learning decision trees in the agnostic model (without membership queries) runs in $n^{O(\log n)}$ [KKMS08] time and with membership queries scales as $\operatorname{poly}(n)$ [GKK08b]. [^3] By membership queries we mean that an algorithm can query an unknown $f:{\{0,1\}}^{n}\rightarrow{\{0,1\}}$ on an $x$ of its choice. In the (uniform) agnostic model, by membership query we mean the following: for every distribution $D:{\{0,1\}}^{n+1}\rightarrow[0,1]$ whose marginal on the first $n$ bits is uniform, the learner can query $x$ and obtain a sample from $b\sim D(x,\cdot)$. The same results hold for juntas since $k$-juntas have size-$2^{k}$ decision trees. In contrast, for DNF formulas, the SOTA algorithm in the agnostic model (without membership queries) is $n^{O(\log n)}$ [KKMS08] and with membership queries is $n^{O(\log\log n)}$ [GKK08a]. Our work shows that one need not use the extra power of membership queries and can achieve an $n^{\log\log n}$-time result with only quantum examples which is better than the state-of-the-art $n^{\log n}$ result using classical examples.
@@
The class of size-$s$ depth-33 circuits is quantum PAC learnable in time $\operatorname{poly}(n,s^{\log\log s},1/\varepsilon)$.
We formally state this theorem in Section 4.4. In contrast, classically, it is a long-standing open problem to efficiently PAC learn even depth-$2$ circuits of size $s$, wherein the state-of-the-art algorithm runs in time $n^{O(\log s)}$. In contrast, Bshouty and Jackson [BJ95] showed depth-$2$ circuits are quantumly learnable in time $\operatorname{poly}(n,s)$ and we further show that even depth-$3$ circuits are learnable in “near-polynomial time,” thereby giving a state-of-the-art separation between quantum PAC and classical PAC learning. We summarize our main contributions in the table below.
@@
Agnostic DNF $n^{O(\log\log n)}$ [GKK08a] $n^{O(\log n)}$ [KMV08] $n^{O(\log\log n)}$ This work
PAC depth-$3$ $n^{O(\log\log n)}$ [Fel09, KK09] $n^{O(\log n)}$ [Fel09, KK09] $n^{O(\log\log n)}$ This work
Table 1: This gives a summary of the state-of-the-art results for agnostic learning $\textsf{AC}^{0}$ circuits of depth-$2$ (i.e., DNFs) and PAC learning depth-$3$ circuits acting on $n$ bits. For simplicity, we assume that the size of the circuits is $\operatorname{poly}(n)$ and the approximation error is $\varepsilon=1/\operatorname{poly}(n)$.
Beyond depth-$3$ circuits, we are also able to learn $\textsf{TAC}_{2}^{0}$, i.e., threshold of depth-$2$ circuits (consisting of and $\and,\textsf{OR},\textsf{NOT}$ gates). This class of circuits is compelling for two reasons. First, even quantum-efficiently learning thresholds of threshold gates would imply breakthrough classical circuit lower bounds (see [AGG+22] for more details). Second, as observed in [AGS21], learning threshold circuits is equivalent to learning weights of feed-forward neural networks. Thus, quantum efficient learnability of threshold circuits would translate into a dramatic advantage over classical computers for training neural networks, echoing Aaronson’s “Ten Semi-Grand Challenges for Quantum Computing Theory” [Aar05]. Prior to our work, no results were known for PAC learning circuits consisting of any threshold gates; ours is the first to handle a single threshold on the top.
1.2 Technical overview
In this section, we give an overview of the proofs of our main results. Before doing so, we first discuss a few potential approaches that do not yield efficient algorithms. These ideas will, in turn, motivate the need for new learning algorithms.
1.2.1 Prior works and conceptual challenges
@@
Although the three approaches above give new agnostic learning algorithms for parities, disjunctions and juntas, as we’ve mentioned these approaches eventually lead to agnostic algorithms for $s$-term DNFs with complexity that is exponential in $s$, and our goal is to ideally have a $\operatorname{poly}(n,s)$ algorithm for this class. Conceptually, the goal of these works is to solve a much harder problem, so they do not immediately yield results for the task that we are concerned with in this work (whose motivation comes more from quantum learning theory). As alluded to earlier, our main contribution is in obtaining an umbrella framework, that achieves two purposes: $(i)$ unifies all the learning algorithms for different classes of phase states, and $(ii)$ is simpler than the algorithms mentioned above (which solve a harder task). We achieve this via quantum agnostic boosting, which we describe in the next section, which could be of independent interest.
@@ 1.2.2 Quantum agnostic boosting
Our quantum agnostic learner is inspired by the gradient-descent based algorithms for classical boosting proposed by the works of Kalai-Kanade [KK09] and Feldman [FGKP09]. Below, we first give a high-level idea of the boosting algorithm before describing the iterations of the procedure.
@@ 1.2.3 Learning algorithms
Agnostic learning DNFs. The agnostic learner for DNFs is similar to the one for decision trees, except that we need to use Mansour’s result [Man92, LT22] that shows that for size-$s$ DNF formulas, the entire Fourier spectrum is concentrated on $s^{O(\log\log s)}$ coefficients. Using a similar argument as the one for deriving Eq. (2), one can show that if $|\psi\rangle$ is $\tau$-close to a DNF phase state, then there exists a basis state $|\chi_{T}\rangle$ such that $|\langle\psi|\chi_{T}\rangle|^{2}\geq\tau/s^{O(\log\log s)}$. Once again using our quantum boosting algorithm, we obtain a $\operatorname{poly}(n,s^{\log\log s},1/\varepsilon)$ algorithm for agnostic learning size-$s$ DNF formulas.
PAC learning depth-$3$ circuits. To learn these circuits, we employ our agnostic DNF learner. The key insight is that when the input state $|\psi_{f}\rangle$ is promised to correspond to a depth-$3$ circuit then state tomography is accomplished when agnostic learning against DNFs is accomplished, which was also observed classically [Fel09]. This follows from using the seminal result of Hajnal et al. [HMP+93] that says that if $f$ is a threshold of $m$ many DNF formulas $\{g_{1},\ldots,g_{m}\}$ (each with size at most $s$), then $|\langle\psi_{f}|\psi_{g_{i}}\rangle|\geq 1/m,$ for some $i\in[m]$, where $|\psi_{f}\rangle$ (in this section) equals $\frac{1}{\sqrt{2^{n}}}\sumop\displaylimits_{x}f(x)|x\rangle$ since we are in the PAC learning setting. We can now use our agnostic DNF learner outputs a quantum state $|\phi\rangle$ which is at least $\textsf{opt}/m^{2}$ close to $f$. This will serve as our weak learner which we will then boost into a strong PAC learner. The boosting algorithm outputs a (classical description of a) quantum state $|\phi\rangle$ which is close to an unknown $|\psi_{f}\rangle$; we now round the final state $|\phi\rangle$ of the algorithm and show that it satisfies the requirement of PAC learning. Since the runtime of the DNF learner scales as $\operatorname{poly}(n,s^{\log\log s})$, the overall quantum PAC learning algorithm scales as $\operatorname{poly}(n,s^{\log\log s},1/\varepsilon)$.
@@
and the goal is to output a function $h:{\{0,1\}}^{n}\rightarrow{\{0,1\}}$ such that $\Pr_{(x,b)\sim D}[h(x)=b]\geq\textsf{opt}-\varepsilon$ where $\textsf{opt}=\max_{c\in{\mathcal{C}}}[c(x)=b]$. In contrast to the first model, in this distributional model the learning algorithm enjoys the benefit of knowing that the unknown quantum state $|\psi\rangle$ has the form of $|\psi_{D}\rangle$. However, the algorithm needs to output a function $h$, whereas in the first model it can output an arbitrary $|\varphi\rangle$. The two situations are therefore incomparable. In Section 5 we show that if $\phi$ is “nice”, i.e., $\mathop{\mathbb{E}}_{x}[\phi(x)^{2}]\leq\textsf{opt}$, having a learning algorithm in the first model implies a learning algorithm in the second model. In particular, for these distributions, we also obtain quantum learning algorithms in the distributional model. We leave the other direction for future work.
@@ 1.3 Outlook
    Learning DNFs: Our agnostic learning algorithm for $s$-term DNFs runs in “near-polynomial time,” $s^{O(\log\log s)}.$ Is there a $\operatorname{poly}(s)$-time quantum algorithm?
@@ Acknowledgments.
SA thanks Matthias Caro, Alex Grilo and Ryan Sweke for an early discussion on agnostic learning. AD thanks Isaac Chuang and Kristan Temme for early discussions on agnostic learning phase states. We thank Varun Kanade for helpful clarifications of the classical SOTA algorithms. This work was done when MdO was an intern at IBM Quantum.
@@ 2.4.1 PAC learning
In particular, the learner is given copies of $|\psi_{c}\rangle$ and is allowed to perform arbitrary measurements on those copies. Note that measuring $|\psi_{c}\rangle$ in the computational basis produces a classical labeled example, so quantum examples are at least as strong as classical examples. Understanding the strengths and weaknesses of quantum examples has been looked at by several works (we refer an interested reader to the survey [ADW17]). As with the classical complexities, one can similarly define the $(\varepsilon,\delta)$-sample and time complexity for learning $\mathcal{C}$} as the quantum sample complexity (i.e., number of quantum examples $|\psi_{c}\rangle$ used) and quantum time complexity (i.e., number of quantum gates used in the algorithm) of an optimal $(\varepsilon,\delta)$-learner for $\mathcal{C}$}.
@@ 2.4.3 Quantum agnostic learning
In this model, the hypothesis class is a set of quantum state $\mathcal{S}=\{|\psi_{1}\rangle,\ldots,|\psi_{m}\rangle\}$. [^12] We remark that we define this model for pure states for simplicity since that is the focus of this work. One could similarly define a hypothesis class of mixed states $\{\rho_{1},\ldots,\rho_{m}\}$. The agnostic learning algorithm is given copies of an unknown $|\phi\rangle$ and the goal is to output an $|\psi^{\prime}\rangle$ such that
@@ 3 Quantum agnostic boosting
Let $\tau\in(0,1)$ and $\eta(\cdot)$ be a function of $\tau$. Suppose $|\phi\rangle$ is an arbitrary $n$-qubit quantum state such that ${\cal F}_{{\mathcal{C}}}(|\phi\rangle)\geq\tau$. We say ${\mathcal{A}}_{\mathsf{WAL}}$ is a weak agnostic learner for ${\cal S}_{\mathcal{C}}$ if, given copies of $|\phi\rangle$, outputs $|\chi\rangle\in{\mathcal{C}}_{\mathsf{Par}}$ such that $|\langle\phi|\chi\rangle|^{2}\geq\eta(\tau)$ with probability $\geq 1-\delta$. Let $S_{\mathsf{WAL}}$ and $T_{\mathsf{WAL}}$ be the sample and time complexity of $\mathcal{A}_{\mathsf{WAL}}$ respectively. [^13] We remark that in this work we only consider weak agnostic learners whose output class will be the class of parity states, hence why we define it this way.
@@
Let ${\mathcal{A}}_{\mathsf{WAL}}$ be a weak agnostic learner for ${\cal S}_{\mathcal{C}}$ with promise of $\eta$ and $S_{\mathsf{WAL}}$ sample complexity along with $T_{\mathsf{WAL}}$ time complexity. Then, there is an algorithm $\mathcal{L}$} that with probability $\geq 1-\delta$, outputs a state $|\widehat{\phi}\rangle$ expressed as

### v2

@@ Abstract
    $s$-term DNF formulas in time $\textsf{poly}(n,(s/\varepsilon)^{\log\log(s/\varepsilon)\cdot\log(1/\varepsilon)})$.
@@
Using quantum agnostic boosting, we obtain a $n^{O(\log\log n\cdot\log(1/\varepsilon))}$-time algorithm for learning $\textsf{poly}(n)$-sized depth-$3$ circuits (consisting of AND, OR, NOT gates) in the uniform PAC model given quantum examples, which is near-polynomial time for constant $\varepsilon$. Classically, obtaining an algorithm with a similar complexity has been an open question in the PAC model and our work answers this given quantum examples.
@@ 1.1 Main results
Our main result is to make progress on both of the questions highlighted above. For both the tasks of agnostically learning $\operatorname{poly}(n)$-sized DNFs and PAC learning $\operatorname{poly}(n)$-sized depth-$3$ circuits, we give $n^{O(\log\log n\cdot\log(1/\varepsilon))}$ time quantum algorithms (which for constant $\varepsilon$ can be viewed as “near-polynomial” time in the PAC model). Below, we will discuss these results, starting with our main technical contribution, quantum agnostic boosting.
@@
Size-$t$ decision trees and $k$-juntas are learnable in time $\operatorname{poly}(n,t,1/\varepsilon)$ and $\operatorname{poly}(n,2^{k},1/\varepsilon)$ respectively. Similarly, $s$-term DNF formulas are learnable in time $\operatorname{poly}(n,(s/\varepsilon)^{\log\log s/\varepsilon\cdot\log(1/\varepsilon)})$.
We formally state this theorem for each class in Section 4. In order to compare with classical results and for ease of exposition, let us consider the parameters above to take values of $2^{k},t,s,1/\varepsilon=\operatorname{poly}(n)$ (which in general are considered the most interesting settings of these parameters for learning). Classically, the SOTA algorithm for learning decision trees in the agnostic model (without membership queries) runs in $n^{O(\log n)}$ [KKMS08] time and with membership queries scales as $\operatorname{poly}(n)$ [GKK08b]. [^3] By membership queries we mean that an algorithm can query an unknown $f:{\{0,1\}}^{n}\rightarrow{\{0,1\}}$ on an $x$ of its choice. In the (uniform) agnostic model, by membership query we mean the following: for every distribution $D:{\{0,1\}}^{n+1}\rightarrow[0,1]$ whose marginal on the first $n$ bits is uniform, the learner can query $x$ and obtain a sample from $b\sim D(x,\cdot)$. The same results hold for juntas since $k$-juntas have size-$2^{k}$ decision trees. In contrast, for DNF formulas, the SOTA algorithm in the agnostic model is $n^{O(\log^{2}n)}$ [KKMS08, GKK08a]. Our work shows that one can achieve an $n^{O(\log n)}$-time result with only quantum examples.
@@
The class of size-$s$ depth-33 circuits is quantum PAC learnable upto error $\varepsilon$ in time $\operatorname{poly}(n,(s/\varepsilon)^{\log\log(s/\varepsilon)\cdot\log(1/\varepsilon)})$.
We formally state this theorem in Section 4.4. Classically, it is a long-standing open problem to efficiently PAC learn depth-$2$ circuits of size $s$, wherein the state-of-the-art algorithm runs in time $n^{O(\log s)}$. In contrast, Bshouty and Jackson [BJ95] showed depth-$2$ circuits are quantumly learnable in time $\operatorname{poly}(n,s)$ and we further show that even depth-$3$ circuits exhibit a separation between quantum PAC and classical PAC learning. Classically the SOTA algorithm for learning $\operatorname{poly}(n)$-sized depth-$3$ circuits (without queries) scales as $n^{O(\log^{2}n)}$ that comes via Fourier concentration bounds of Tal [Tal17]. We summarize our main contributions in the table below.
@@
Agnostic DNF
$n^{\log(1/\varepsilon)\log\log n}$
[GKK08a]
$n^{\log(1/\varepsilon)\log n}$
[KKMS08]
$n^{\log(1/\varepsilon)\log\log n}$
This work
PAC depth-$3$
$n^{\log(1/\varepsilon)\log\log n}$
[Fel09, KK09]
$n^{\log(1/\varepsilon)\log n}$
[Tal17]
$n^{\log(1/\varepsilon)\log\log n}$
This work
Table 1: This gives a summary of the state-of-the-art results for agnostic learning $\operatorname{poly}(n)$-sized $\textsf{AC}^{0}$ circuits of depth-$2$ (i.e., DNFs) and PAC learning depth-$3$ circuits acting on $n$ bits.
Apart from depth-$3$ circuits, we are able to learn $\textsf{TAC}_{2}^{0}$, i.e., threshold of depth-$2$ circuits (consisting of and $\and,\textsf{OR},\textsf{NOT}$ gates). This class of circuits is compelling for two reasons. First, even quantum-efficiently learning thresholds of threshold gates would imply breakthrough classical circuit lower bounds (see [AGG+22] for more details). Second, as observed in [AGS21], learning threshold circuits is equivalent to learning weights of feed-forward neural networks. Thus, quantum efficient learnability of threshold circuits would translate into a dramatic advantage over classical computers for training neural networks, echoing Aaronson’s “Ten Semi-Grand Challenges for Quantum Computing Theory” [Aar05]. Prior to our work, no results were known for PAC learning circuits consisting of any threshold gates; ours is the first to handle a single threshold on the top.
1.2 Prior works and concept challenges
We first discuss a few potential approaches that do not yield efficient algorithms. These ideas will, in turn, motivate the need for new learning algorithms.
1.2.1 In the quantum setting
@@
1.2.2 In the classical setting
In classical learning theory, agnostic learning decision trees (DT) and DNFs has received a lot of attention. Classically, the model of agnostic learning is defined as follows: a learner is given uniformly random $x\in{\{0,1\}}^{n}$ and $b\sim(1+\phi(x)/2,1-\phi(x)/2)$ and the goal is to find a function $f$ from the concept class which agrees with $\phi$ as well as possible. The SOTA algorithms for size-$s$ DTs run in time polynomial in $n,s$ but requires membership queries [Fel09, GKK08b]. They crucially use that the $\ell_{1}$ norm of the Fourier coefficients of $f\in\textsf{DT}(s)$ is at most $s$. However, DNFs do not enjoy the property of small $\ell_{1}$ norm, but instead are only known to have sparse Fourier spectrum and at this point, it is even unclear classically how one can obtain a weak learner for DNFs (with examples) if $\phi$ is only promised to satisfy $\max_{x}|\phi(x)|\leq 1$. We show that this can be circumvented when given access to quantum examples and a weak agnostic learner can be naturally proposed with just the Fourier concentration promise.
@@ 1.2.2 In the classical setting
Although the three quantum approaches give new agnostic learning algorithms for parities, disjunctions and juntas, as we’ve mentioned these approaches eventually lead to agnostic algorithms for $s$-term DNFs with complexity that is exponential in $s$, and our goal is to ideally have a $\operatorname{poly}(n,s)$ algorithm for this class. Conceptually, the goal of these works is to solve a much harder problem, so they do not immediately yield results for the task that we are concerned with in this work (whose motivation comes more from quantum learning theory). As alluded to earlier, our main contribution is in obtaining an umbrella framework, that achieves two purposes: $(i)$ unifies all the learning algorithms for different classes of phase states, and $(ii)$ is simpler than the algorithms mentioned above (which solve a harder task). We achieve this via quantum agnostic boosting, which we describe in the next section, which could be of independent interest.
1.3 Technical overview
In this section, we give an overview of the proofs of our main results.
@@ 1.3.1 Quantum agnostic boosting
Our quantum agnostic learner is inspired by the gradient-descent based algorithms for classical boosting proposed by the works of Kalai-Kanade [KK09] and Feldman [Fel09]. Below, we first give a high-level idea of the boosting algorithm before describing the iterations of the procedure.
@@ 1.3.2 Learning algorithms
Agnostic learning DNFs. The agnostic learner for DNFs is similar to the one for decision trees, except that we need to use Mansour’s result [Man92, LT22] that shows that for size-$s$ DNF formulas, the entire Fourier spectrum is concentrated on $s^{O(\log\log s\log(1/\varepsilon))}$ coefficients. Using a similar argument as the one for deriving Eq. (2), one can show that if $|\psi\rangle$ is $\tau$-close to a DNF phase state, then there exists a basis state $|\chi_{T}\rangle$ such that $|\langle\psi|\chi_{T}\rangle|^{2}\geq\tau/s^{O(\log\log s\log(1/\varepsilon))}$. Once again using our quantum boosting algorithm, we obtain a $\operatorname{poly}(n,s^{\log\log s\log(1/\varepsilon)})$ algorithm for agnostic learning size-$s$ DNF formulas.
PAC learning depth-$3$ circuits. To learn these circuits, we employ our agnostic DNF learner. The key insight is that when the input state $|\psi_{f}\rangle$ is promised to correspond to a depth-$3$ circuit then state tomography is accomplished when agnostic learning against DNFs is accomplished, which was also observed classically [Fel09]. This follows from using the seminal result of Hajnal et al. [HMP+93] that says that if $f$ is a threshold of $m$ many DNF formulas $\{g_{1},\ldots,g_{m}\}$ (each with size at most $s$), then $|\langle\psi_{f}|\psi_{g_{i}}\rangle|\geq 1/m,$ for some $i\in[m]$, where $|\psi_{f}\rangle$ (in this section) equals $\frac{1}{\sqrt{2^{n}}}\sumop\displaylimits_{x}f(x)|x\rangle$ since we are in the PAC learning setting. We can now use our agnostic DNF learner outputs a quantum state $|\phi\rangle$ which is at least $\textsf{opt}/m^{2}$ close to $f$. This will serve as our weak learner which we will then boost into a strong PAC learner. The boosting algorithm outputs a (classical description of a) quantum state $|\phi\rangle$ which is close to an unknown $|\psi_{f}\rangle$; we now round the final state $|\phi\rangle$ of the algorithm and show that it satisfies the requirement of PAC learning. Since the runtime of the DNF learner scales as $\operatorname{poly}(n,s^{\log\log s\log(1/\varepsilon)})$, the overall quantum PAC learning algorithm scales as $\operatorname{poly}(n,s^{\log\log s\log(1/\varepsilon)})$.
@@
and the goal is to output a function $h:{\{0,1\}}^{n}\rightarrow{\{0,1\}}$ such that $\Pr_{(x,b)\sim D}[h(x)=b]\geq\textsf{opt}-\varepsilon$ where $\textsf{opt}=\max_{c\in{\mathcal{C}}}[c(x)=b]$. In contrast to the first model, in this distributional model the learning algorithm enjoys the benefit of knowing that the unknown quantum state $|\psi\rangle$ has the form of $|\psi_{D}\rangle$. However, the algorithm needs to output a function $h$, whereas in the first model it can output an arbitrary $|\varphi\rangle$. The two situations are therefore incomparable and interesting for different reasons. In Section 5 we show that if $\phi$ is “nice”, i.e., $\mathop{\mathbb{E}}_{x}[\phi(x)^{2}]\leq\textsf{opt}$, having a learning algorithm in the first model implies a learning algorithm in the second model. In particular, for these distributions, we also obtain quantum learning algorithms in the distributional model. We leave the other direction for future work.
@@ 1.4 Outlook
    Learning depth-$3$ circuits: Our PAC learning algorithm for depth-$3$ circuits (with constant $\varepsilon$) runs in “near-polynomial time,” $n^{O(\log\log n)}.$ Is there a $\operatorname{poly}(n)$-time quantum algorithm?
@@ Acknowledgments.
SA thanks Matthias Caro, Alex Grilo and Ryan Sweke for an early discussion on agnostic learning. AD thanks Isaac Chuang and Kristan Temme for early discussions on agnostic learning phase states. We thank Igor Carboni, Gautam Chandrasekaran, Varun Kanade, and Adam Klivans for helpful clarifications on the classical SOTA algorithms. This work was done when MdO was an intern at IBM Quantum.
@@ 2.4.1 PAC learning
In particular, the learner is given copies of $|\psi_{c}\rangle$ and is allowed to perform arbitrary measurements on those copies. Note that measuring $|\psi_{c}\rangle$ in the computational basis produces a classical labeled example, so quantum examples are at least as strong as classical examples. Understanding the strengths and weaknesses of quantum examples has been looked at by several works (we refer an interested reader to the survey [ADW17]). As with the classical complexities, one can similarly define the $(\varepsilon,\delta)$-sample and time complexity for learning $\mathcal{C}$} as the quantum sample complexity (i.e., number of quantum examples $|\psi_{c}\rangle$ used) and quantum time complexity (i.e., number of one and two-qubit quantum gates used in the algorithm) of an optimal $(\varepsilon,\delta)$-learner for $\mathcal{C}$}.
@@ 2.4.3 Quantum agnostic learning
In this model, the hypothesis class is a set of quantum state $\mathcal{S}=\{|\psi_{1}\rangle,\ldots,|\psi_{m}\rangle\}$. [^12] We remark that we define this model for pure states for simplicity since that is the focus of this work. One could similarly define a hypothesis class of mixed states $\{\rho_{1},\ldots,\rho_{m}\}$. The agnostic learning algorithm is given copies of an unknown $|\phi\rangle$ and the goal is to output an $|\phi^{\prime}\rangle$ such that
@@ 3 Quantum agnostic boosting
Let $\tau\in(0,1)$ and $\eta(\cdot)$ be a function of $\tau$. Suppose $|\phi\rangle$ is an arbitrary $n$-qubit quantum state such that ${\cal F}_{{\mathcal{C}}}(|\phi\rangle)\geq\tau$. We say ${\mathcal{A}}_{\mathsf{WAL}}$ is a weak agnostic learner for ${\cal S}_{\mathcal{C}}$ with promise $\eta(\cdot)$ if, given copies of $|\phi\rangle$, outputs $|\chi\rangle\in{\mathcal{C}}_{\mathsf{Par}}$ such that $|\langle\phi|\chi\rangle|^{2}\geq\eta(\tau)$ with probability $\geq 1-\delta$. Let $S_{\mathsf{WAL}}$ and $T_{\mathsf{WAL}}$ be the sample and time complexity of $\mathcal{A}_{\mathsf{WAL}}$ respectively. [^13] We remark that in this work we only consider weak agnostic learners whose output class will be the class of parity states, hence why we define it this way.
@@
Let ${\mathcal{A}}_{\mathsf{WAL}}$ be a weak agnostic learner for ${\cal S}_{\mathcal{C}}$ with promise of $\eta(\cdot)$, with sample complexity $S_{\mathsf{WAL}}$ and time complexity $T_{\mathsf{WAL}}$. Then, there is an algorithm $\mathcal{L}$} that with probability $\geq 1-\delta$, outputs a state $|\widehat{\phi}\rangle$ expressed as